# Bangkok 10330, Thailand

### v1

@@ 3.1 Results and time independence
Here, we review the bispectrum up to NLO Antoniadis:2024abm in single field inflation and discuss its time independence. We first introduce the amplitude $\mathscr{A}$ of the bispectrum $\mathcal{B}$ through the Hubble and slow-roll parameters at the time $\tau$ when the 3-point function is defined:
$\mathcal{B}(k_{1},k_{2},k_{3})=\frac{H^{4}}{32\varepsilon_{1}^{2}}\frac{\mathscr{A}}{k_{1}^{3}k_{2}^{3}k_{3}^{3}}.$ (16)
@@
$\mathscr{A}$ $=\mathscr{A}_{\text{LO}}+\mathscr{A}_{\text{NLO}}+\mathcal{O}(\varepsilon^{3}),$ (17)
@@
$\mathscr{A}_{\text{LO}}$ $=\varepsilon_{1}K_{1}\big{(}k_{1},k_{2},k_{3}\big{)}+\varepsilon_{2}K_{2}\big{(}k_{1},k_{2},k_{3}\big{)}+\mathcal{O}(\tau),$ (18)
$\begin{split}\mathscr{A}_{\text{NLO}}&=\left[\gamma_{E}-1+\ln(\!-\tau k_{\text{t}})\right]\Big{(}\varepsilon_{1}^{2}K_{3}\big{(}k_{1},k_{2},k_{3}\big{)}\\ &\qquad\quad+\varepsilon_{1}\varepsilon_{2}K_{4}\big{(}k_{1},k_{2},k_{3}\big{)}+\varepsilon_{2}^{2}K_{5}\big{(}k_{1},k_{2},k_{3}\big{)}+\varepsilon_{2}\varepsilon_{3}K_{6}\big{(}k_{1},k_{2},k_{3}\big{)}\Big{)}\\ &\quad~{}+\big{(}2\varepsilon_{1}+\varepsilon_{2}\big{)}\ln\left(\frac{2k_{1}}{k_{\text{t}}}\right)\Big{(}\varepsilon_{1}K_{7}\big{(}k_{1},k_{2},k_{3}\big{)}+\varepsilon_{2}K_{8}\big{(}k_{1},k_{2},k_{3}\big{)}\Big{)}\\ &\quad~{}+\big{(}2\varepsilon_{1}+\varepsilon_{2}\big{)}\ln\left(\frac{2k_{2}}{k_{\text{t}}}\right)\Big{(}\varepsilon_{1}K_{7}\big{(}k_{2},k_{1},k_{3}\big{)}+\varepsilon_{2}K_{8}\big{(}k_{2},k_{1},k_{3}\big{)}\Big{)}\\ &\quad~{}+\big{(}2\varepsilon_{1}+\varepsilon_{2}\big{)}\ln\left(\frac{2k_{3}}{k_{\text{t}}}\right)\Big{(}\varepsilon_{1}K_{7}\big{(}k_{3},k_{2},k_{1}\big{)}+\varepsilon_{2}K_{8}\big{(}k_{3},k_{2},k_{1}\big{)}\Big{)}\\ &\quad~{}+\varepsilon_{1}^{2}K_{9}\big{(}k_{1},k_{2},k_{3}\big{)}+\varepsilon_{1}\varepsilon_{2}K_{10}\big{(}k_{1},k_{2},k_{3}\big{)}\\ &\quad~{}+\varepsilon_{2}^{2}K_{11}\big{(}k_{1},k_{2},k_{3}\big{)}+\varepsilon_{2}\varepsilon_{3}K_{12}\big{(}k_{1},k_{2},k_{3}\big{)}+\mathcal{O}(\tau),\end{split}$ (19)
@@
$\begin{split}K_{1}\big{(}k_{1},k_{2},k_{3}\big{)}&=8\frac{k_{1}^{2}k_{2}^{2}+k_{1}^{2}k_{3}^{2}+k_{2}^{2}k_{3}^{2}}{k_{\text{t}}}-2(k_{1}^{3}+k_{2}^{3}+k_{3}^{3})+k_{\text{t}}(k_{1}^{2}+k_{2}^{2}+k_{3}^{2}),\\ K_{2}\big{(}k_{1},k_{2},k_{3}\big{)}&=k_{1}^{3}+k_{2}^{3}+k_{3}^{3},\\ K_{3}\big{(}k_{1},k_{2},k_{3}\big{)}&=-4K_{1}\big{(}k_{1},k_{2},k_{3}\big{)},\\ K_{4}\big{(}k_{1},k_{2},k_{3}\big{)}&=-K_{1}\big{(}k_{1},k_{2},k_{3}\big{)}-4K_{2}\big{(}k_{1},k_{2},k_{3}\big{)},\\ K_{5}\big{(}k_{1},k_{2},k_{3}\big{)}&=-2K_{2}\big{(}k_{1},k_{2},k_{3}\big{)},\\ K_{6}\big{(}k_{1},k_{2},k_{3}\big{)}&=K_{2}\big{(}k_{1},k_{2},k_{3}\big{)},\\ K_{7}\big{(}k_{1},k_{2},k_{3}\big{)}&=\frac{k_{2}+k_{3}}{-k_{1}+k_{2}+k_{3}}\Big{(}2k_{1}k_{2}k_{3}-K_{1}\big{(}k_{1},k_{2},k_{3}\big{)}\Big{)},\\ K_{8}\big{(}k_{1},k_{2},k_{3}\big{)}&=k_{1}^{3}-K_{2}\big{(}k_{1},k_{2},k_{3}\big{)},\\ K_{9}\big{(}k_{1},k_{2},k_{3}\big{)}&=2k_{1}k_{2}k_{3}-K_{1}\big{(}k_{1},k_{2},k_{3}\big{)},\\ K_{10}\big{(}k_{1},k_{2},k_{3}\big{)}&=2K_{1}\big{(}k_{1},k_{2},k_{3}\big{)}-3K_{2}\big{(}k_{1},k_{2},k_{3}\big{)}-3k_{1}k_{2}k_{3}+k_{\text{t}}(k_{1}^{2}+k_{2}^{2}+k_{3}^{2}),\\ K_{11}\big{(}k_{1},k_{2},k_{3}\big{)}&=2K_{2}\big{(}k_{1},k_{2},k_{3}\big{)},\\ K_{12}\big{(}k_{1},k_{2},k_{3}\big{)}&=k_{1}k_{2}k_{3}+K_{2}\big{(}k_{1},k_{2},k_{3}\big{)}-k_{\text{t}}(k_{1}^{2}+k_{2}^{2}+k_{3}^{2}).\end{split}$ (20)
@@
Let us consider the time (in)dependence of the bispectrum. In general, the computation of higher order corrections in slow-roll parameters involves the expansion of the Hubble and slow-roll parameters in terms of their small time derivatives (namely, the slow-roll parameters) around a reference time, which we call a pivot time. It can be chosen arbitrarily. For example, in the computation of the bispectrum in Antoniadis:2024abm , the pivot time was chosen to be the time of the correlator $\tau$, so that the Hubble and slow-roll parameters in $\mathcal{B}$ above are evaluated at $\tau$. Therefore, the dependence of $\mathcal{B}$ on $\tau$ comes from not only the logarithmic term $\ln(-k_{\text{t}}\tau)$ but also the parameters $H,\varepsilon_{i}$. A natural question is then how the expression of the bispectrum $\mathcal{B}$ changes when $H,\varepsilon_{i}$ are expanded around another pivot time $\tau^{*}$. After the expansion, the bispectrum is expressed in $H,\varepsilon_{i}$ evaluated at $\tau^{*}$ and will contain new powers of $\ln(\tau/\tau^{*})$ which come from the expansion [^5] This logarithm is nothing but $H(t_{*}-t)$ in terms of the proper time at leading order in the slow-roll parameters.. It should be noted that this is just rewriting the bispectrum, manifesting the $\tau$ dependence of $H,\varepsilon_{i}$ through their expansion. Now, if the bispectrum after the expansion around $\tau^{*}$ turns out to be independent of $\tau$, one may claim that the bispectrum, which is defined at time $\tau$, is actually independent of $\tau$. To put it simply, this means that the $\tau$ dependence of $H,\varepsilon_{i}$ compensates the pre-existing logarithm $\ln(-\tau)$.
@@
$\mathcal{B}_{\text{LO}}=\frac{H^{4}}{32\varepsilon_{1}^{2}}\frac{\mathscr{A}_{\text{LO}}}{k_{1}^{3}k_{2}^{3}k_{3}^{3}},\quad\mathcal{B}_{\text{NLO}}=\frac{H^{4}}{32\varepsilon_{1}^{2}}\frac{\mathscr{A}_{\text{NLO}}}{k_{1}^{3}k_{2}^{3}k_{3}^{3}},$ (21)
@@
$H(t(\tau))$ $=H^{*}\left[1+\varepsilon_{1}^{*}\ln\left(\frac{\tau}{\tau^{*}}\right)+\mathcal{O}(\varepsilon^{*2})\right],$ (22)
$\varepsilon_{i}(t(\tau))$ $=\varepsilon_{i}^{*}\left[1-\varepsilon_{i+1}^{*}\ln\left(\frac{\tau}{\tau^{*}}\right)+\mathcal{O}(\varepsilon^{*2})\right],$ (23)
@@
$\quad+\ln\left(\frac{\tau}{\tau^{*}}\right)\big{[}\varepsilon_{1}^{*}(4\varepsilon_{1}^{*}+\varepsilon_{2}^{*})K_{1}(k_{1},k_{2},k_{3})$
$\qquad\quad+\varepsilon_{2}^{*}(4\varepsilon_{1}^{*}+2\varepsilon_{2}^{*}-\varepsilon_{3}^{*})K_{2}(k_{1},k_{2},k_{3})\big{]}+\mathcal{O}(\varepsilon^{*3},\tau)\Big{)}$
$\begin{split}&=\frac{H^{*4}}{32\varepsilon_{1}^{*2}}\frac{1}{k_{1}^{3}k_{2}^{3}k_{3}^{3}}\Big{(}\varepsilon_{1}^{*}K_{1}(k_{1},k_{2},k_{3})+\varepsilon_{2}^{*}K_{2}(k_{1},k_{2},k_{3})\\ &\quad-\ln\left(\frac{\tau}{\tau^{*}}\right)\big{[}\varepsilon_{1}^{*2}K_{3}(k_{1},k_{2},k_{3})+\varepsilon_{1}^{*}\varepsilon_{2}^{*}K_{4}(k_{1},k_{2},k_{3})\\ &\qquad+\varepsilon_{2}^{*2}K_{5}(k_{1},k_{2},k_{3})+\varepsilon_{2}^{*}\varepsilon_{3}^{*}K_{6}(k_{1},k_{2},k_{3})\big{]}+\mathcal{O}\big{(}\varepsilon^{*3},\tau\big{)}\Big{)}.\end{split}$ (24)
@@
$\mathcal{B}_{\text{NLO}}(k_{1},k_{2},k_{3})=\frac{H^{4}}{32\varepsilon_{1}^{2}}\frac{1}{k_{1}^{3}k_{2}^{3}k_{3}^{3}}\mathscr{A}_{\text{NLO}}=\frac{H^{*4}}{32\varepsilon_{1}^{*2}}\frac{1}{k_{1}^{3}k_{2}^{3}k_{3}^{3}}\Big{(}\mathscr{A}_{\text{NLO}}^{*}+\mathcal{O}\big{(}\varepsilon^{*3}\big{)}\Big{)},$ (25)
where $\mathscr{A}_{\text{NLO}}^{*}$ is simply $\mathscr{A}_{\text{NLO}}$ with all $\varepsilon_{i}$ replaced by those evaluated at $\tau^{*}$ but $\ln(-k_{\text{t}}\tau)$ untouched. However, in their sum $\mathcal{B}_{\text{LO}}+\mathcal{B}_{\text{NLO}}$, the logarithms $\ln(-\tau)$ cancel out, and the total bispectrum $\mathcal{B}$ turns out to be equal to the expression (16) with the replacements: $\varepsilon_{i}\rightarrow\varepsilon_{i}^{*}$, $H\rightarrow H^{*}$ and $\tau\rightarrow\tau^{*}$:
$\mathcal{B}(k_{1},k_{2},k_{3})=\frac{H^{*4}}{32\varepsilon_{1}^{*2}}\frac{\mathscr{A}^{*}}{k_{1}^{3}k_{2}^{3}k_{3}^{3}},$ (26)
$\mathscr{A}^{*}=(\mathscr{A}_{\text{LO}}+\mathscr{A}_{\text{NLO}})_{\varepsilon_{i}\rightarrow\varepsilon_{i}^{*},\,H\rightarrow H^{*},\,\tau\rightarrow\tau^{*}}+\mathcal{O}(\varepsilon_{i}^{*3}).$ (27)
Therefore, the bispectrum, which is defined at late time $\tau$, is independent of $\tau$ up to NLO as a consequence of the cancellation of $\ln(-\tau)$, meaning that the $\tau$-dependence of $H,\varepsilon_{i}$ in (16) with (17) indeed cancels the logarithm $\ln(-\tau)$ in $\mathscr{A}_{\text{NLO}}$. Furthermore, one can confirm that the bispectrum (26) is independent of the choice of the pivot time $\tau^{*}$ by expanding $H^{*},\varepsilon_{i}^{*}$ around yet another pivot time $\tau^{\circ}$, which gives (26) with the replacement $(H^{*},\varepsilon_{i}^{*},\tau^{*})\to(H^{\circ},\varepsilon_{i}^{\circ},\tau^{\circ})$.
3.2 Dominant contributions to non-gaussianity and its covariant definition

### v2

@@ 3.1 Results and time independence
Here, we review the bispectrum up to NLO [4] in single field inflation and discuss its time independence. We first introduce the amplitude $\mathscr{A}$ of the bispectrum $\mathcal{B}$ through the Hubble and slow-roll parameters at the time $\tau$ when the 3-point function is defined:
$\mathcal{B}(k_{1},k_{2},k_{3})=\frac{H^{4}}{32\varepsilon_{1}^{2}}\frac{\mathscr{A}}{k_{1}^{3}k_{2}^{3}k_{3}^{3}}.$ (3.1)
@@
$\mathscr{A}$ $=\mathscr{A}_{\text{LO}}+\mathscr{A}_{\text{NLO}}+\mathcal{O}(\varepsilon^{3}),$ (3.2)
@@
$\mathscr{A}_{\text{LO}}$ $=\varepsilon_{1}K_{1}\big{(}k_{1},k_{2},k_{3}\big{)}+\varepsilon_{2}K_{2}\big{(}k_{1},k_{2},k_{3}\big{)}+\mathcal{O}(\tau),$ (3.3)
$\begin{split}\mathscr{A}_{\text{NLO}}&=\left[\gamma_{E}-1+\ln(\!-\tau k_{\text{t}})\right]\Big{(}\varepsilon_{1}^{2}K_{3}\big{(}k_{1},k_{2},k_{3}\big{)}\\ &\qquad\quad+\varepsilon_{1}\varepsilon_{2}K_{4}\big{(}k_{1},k_{2},k_{3}\big{)}+\varepsilon_{2}^{2}K_{5}\big{(}k_{1},k_{2},k_{3}\big{)}+\varepsilon_{2}\varepsilon_{3}K_{6}\big{(}k_{1},k_{2},k_{3}\big{)}\Big{)}\\ &\quad~{}+\big{(}2\varepsilon_{1}+\varepsilon_{2}\big{)}\ln\left(\frac{2k_{1}}{k_{\text{t}}}\right)\Big{(}\varepsilon_{1}K_{7}\big{(}k_{1},k_{2},k_{3}\big{)}+\varepsilon_{2}K_{8}\big{(}k_{1},k_{2},k_{3}\big{)}\Big{)}\\ &\quad~{}+\big{(}2\varepsilon_{1}+\varepsilon_{2}\big{)}\ln\left(\frac{2k_{2}}{k_{\text{t}}}\right)\Big{(}\varepsilon_{1}K_{7}\big{(}k_{2},k_{1},k_{3}\big{)}+\varepsilon_{2}K_{8}\big{(}k_{2},k_{1},k_{3}\big{)}\Big{)}\\ &\quad~{}+\big{(}2\varepsilon_{1}+\varepsilon_{2}\big{)}\ln\left(\frac{2k_{3}}{k_{\text{t}}}\right)\Big{(}\varepsilon_{1}K_{7}\big{(}k_{3},k_{2},k_{1}\big{)}+\varepsilon_{2}K_{8}\big{(}k_{3},k_{2},k_{1}\big{)}\Big{)}\\ &\quad~{}+\varepsilon_{1}^{2}K_{9}\big{(}k_{1},k_{2},k_{3}\big{)}+\varepsilon_{1}\varepsilon_{2}K_{10}\big{(}k_{1},k_{2},k_{3}\big{)}\\ &\quad~{}+\varepsilon_{2}^{2}K_{11}\big{(}k_{1},k_{2},k_{3}\big{)}+\varepsilon_{2}\varepsilon_{3}K_{12}\big{(}k_{1},k_{2},k_{3}\big{)}+\mathcal{O}(\tau),\end{split}$ (3.4)
@@
$\begin{split}K_{1}\big{(}k_{1},k_{2},k_{3}\big{)}&=8\frac{k_{1}^{2}k_{2}^{2}+k_{1}^{2}k_{3}^{2}+k_{2}^{2}k_{3}^{2}}{k_{\text{t}}}-2(k_{1}^{3}+k_{2}^{3}+k_{3}^{3})+k_{\text{t}}(k_{1}^{2}+k_{2}^{2}+k_{3}^{2}),\\ K_{2}\big{(}k_{1},k_{2},k_{3}\big{)}&=k_{1}^{3}+k_{2}^{3}+k_{3}^{3},\\ K_{3}\big{(}k_{1},k_{2},k_{3}\big{)}&=-4K_{1}\big{(}k_{1},k_{2},k_{3}\big{)},\\ K_{4}\big{(}k_{1},k_{2},k_{3}\big{)}&=-K_{1}\big{(}k_{1},k_{2},k_{3}\big{)}-4K_{2}\big{(}k_{1},k_{2},k_{3}\big{)},\\ K_{5}\big{(}k_{1},k_{2},k_{3}\big{)}&=-2K_{2}\big{(}k_{1},k_{2},k_{3}\big{)},\\ K_{6}\big{(}k_{1},k_{2},k_{3}\big{)}&=K_{2}\big{(}k_{1},k_{2},k_{3}\big{)},\\ K_{7}\big{(}k_{1},k_{2},k_{3}\big{)}&=\frac{k_{2}+k_{3}}{-k_{1}+k_{2}+k_{3}}\Big{(}2k_{1}k_{2}k_{3}-K_{1}\big{(}k_{1},k_{2},k_{3}\big{)}\Big{)},\\ K_{8}\big{(}k_{1},k_{2},k_{3}\big{)}&=k_{1}^{3}-K_{2}\big{(}k_{1},k_{2},k_{3}\big{)},\\ K_{9}\big{(}k_{1},k_{2},k_{3}\big{)}&=2k_{1}k_{2}k_{3}-K_{1}\big{(}k_{1},k_{2},k_{3}\big{)},\\ K_{10}\big{(}k_{1},k_{2},k_{3}\big{)}&=2K_{1}\big{(}k_{1},k_{2},k_{3}\big{)}-3K_{2}\big{(}k_{1},k_{2},k_{3}\big{)}-3k_{1}k_{2}k_{3}+k_{\text{t}}(k_{1}^{2}+k_{2}^{2}+k_{3}^{2}),\\ K_{11}\big{(}k_{1},k_{2},k_{3}\big{)}&=2K_{2}\big{(}k_{1},k_{2},k_{3}\big{)},\\ K_{12}\big{(}k_{1},k_{2},k_{3}\big{)}&=k_{1}k_{2}k_{3}+K_{2}\big{(}k_{1},k_{2},k_{3}\big{)}-k_{\text{t}}(k_{1}^{2}+k_{2}^{2}+k_{3}^{2}).\end{split}$ (3.5)
@@
Let us consider the time (in)dependence of the bispectrum. In general, the computation of higher order corrections in slow-roll parameters involves the expansion of the Hubble and slow-roll parameters in terms of their small time derivatives (namely, the slow-roll parameters) around a reference time, which we call a pivot time. It can be chosen arbitrarily. For example, in the computation of the bispectrum in [4], the pivot time was chosen to be the time of the correlator $\tau$, so that the Hubble and slow-roll parameters in $\mathcal{B}$ above are evaluated at $\tau$. Therefore, the dependence of $\mathcal{B}$ on $\tau$ comes from not only the logarithmic term $\ln(-k_{\text{t}}\tau)$ but also the parameters $H,\varepsilon_{i}$. A natural question is then how the expression of the bispectrum $\mathcal{B}$ changes when $H,\varepsilon_{i}$ are expanded around another pivot time $\tau^{*}$. After the expansion, the bispectrum is expressed in $H,\varepsilon_{i}$ evaluated at $\tau^{*}$ and will contain new powers of $\ln(\tau/\tau^{*})$ which come from the expansion [^5] This logarithm is nothing but $H(t_{*}-t)$ in terms of the proper time at leading order in the slow-roll parameters.. It should be noted that this is just rewriting the bispectrum, manifesting the $\tau$ dependence of $H,\varepsilon_{i}$ through their expansion. Now, if the bispectrum after the expansion around $\tau^{*}$ turns out to be independent of $\tau$, one may claim that the bispectrum, which is defined at time $\tau$, is actually independent of $\tau$. To put it simply, this means that the $\tau$ dependence of $H,\varepsilon_{i}$ compensates the pre-existing logarithm $\ln(-\tau)$.
@@
$\mathcal{B}_{\text{LO}}=\frac{H^{4}}{32\varepsilon_{1}^{2}}\frac{\mathscr{A}_{\text{LO}}}{k_{1}^{3}k_{2}^{3}k_{3}^{3}},\quad\mathcal{B}_{\text{NLO}}=\frac{H^{4}}{32\varepsilon_{1}^{2}}\frac{\mathscr{A}_{\text{NLO}}}{k_{1}^{3}k_{2}^{3}k_{3}^{3}},$ (3.6)
@@
$H(t(\tau))$ $=H^{*}\left[1+\varepsilon_{1}^{*}\ln\left(\frac{\tau}{\tau^{*}}\right)+\mathcal{O}(\varepsilon^{*2})\right],$ (3.7)
$\varepsilon_{i}(t(\tau))$ $=\varepsilon_{i}^{*}\left[1-\varepsilon_{i+1}^{*}\ln\left(\frac{\tau}{\tau^{*}}\right)+\mathcal{O}(\varepsilon^{*2})\right],$ (3.8)
@@
$\qquad+\ln\left(\frac{\tau}{\tau^{*}}\right)\big{[}\varepsilon_{1}^{*}(4\varepsilon_{1}^{*}+\varepsilon_{2}^{*})K_{1}(k_{1},k_{2},k_{3})$
$\qquad\qquad\qquad\quad+\varepsilon_{2}^{*}(4\varepsilon_{1}^{*}+2\varepsilon_{2}^{*}-\varepsilon_{3}^{*})K_{2}(k_{1},k_{2},k_{3})\big{]}+\mathcal{O}(\varepsilon^{*3},\tau)\Big{)}$
$\begin{split}&=\frac{H^{*4}}{32\varepsilon_{1}^{*2}}\frac{1}{k_{1}^{3}k_{2}^{3}k_{3}^{3}}\Big{(}\varepsilon_{1}^{*}K_{1}(k_{1},k_{2},k_{3})+\varepsilon_{2}^{*}K_{2}(k_{1},k_{2},k_{3})\\ &\qquad-\ln\left(\frac{\tau}{\tau^{*}}\right)\big{[}\varepsilon_{1}^{*2}K_{3}(k_{1},k_{2},k_{3})+\varepsilon_{1}^{*}\varepsilon_{2}^{*}K_{4}(k_{1},k_{2},k_{3})\\ &\qquad\qquad\qquad\quad+\varepsilon_{2}^{*2}K_{5}(k_{1},k_{2},k_{3})+\varepsilon_{2}^{*}\varepsilon_{3}^{*}K_{6}(k_{1},k_{2},k_{3})\big{]}+\mathcal{O}\big{(}\varepsilon^{*3},\tau\big{)}\Big{)}.\end{split}$ (3.9)
@@
$\mathcal{B}_{\text{NLO}}(k_{1},k_{2},k_{3})=\frac{H^{4}}{32\varepsilon_{1}^{2}}\frac{1}{k_{1}^{3}k_{2}^{3}k_{3}^{3}}\mathscr{A}_{\text{NLO}}=\frac{H^{*4}}{32\varepsilon_{1}^{*2}}\frac{1}{k_{1}^{3}k_{2}^{3}k_{3}^{3}}\Big{(}\mathscr{A}_{\text{NLO}}^{*}+\mathcal{O}\big{(}\varepsilon^{*3}\big{)}\Big{)},$ (3.10)
where $\mathscr{A}_{\text{NLO}}^{*}$ is simply $\mathscr{A}_{\text{NLO}}$ with all $\varepsilon_{i}$ replaced by those evaluated at $\tau^{*}$ but $\ln(-k_{\text{t}}\tau)$ untouched. However, in their sum $\mathcal{B}_{\text{LO}}+\mathcal{B}_{\text{NLO}}$, the logarithms $\ln(-\tau)$ cancel out, and the total bispectrum $\mathcal{B}$ turns out to be equal to the expression (3.1) with the replacements: $\varepsilon_{i}\rightarrow\varepsilon_{i}^{*}$, $H\rightarrow H^{*}$ and $\tau\rightarrow\tau^{*}$:
$\mathcal{B}(k_{1},k_{2},k_{3})=\frac{H^{*4}}{32\varepsilon_{1}^{*2}}\frac{\mathscr{A}^{*}}{k_{1}^{3}k_{2}^{3}k_{3}^{3}},$ (3.11)
$\mathscr{A}^{*}=(\mathscr{A}_{\text{LO}}+\mathscr{A}_{\text{NLO}})_{\varepsilon_{i}\rightarrow\varepsilon_{i}^{*},\,H\rightarrow H^{*},\,\tau\rightarrow\tau^{*}}+\mathcal{O}(\varepsilon_{i}^{*3}).$ (3.12)
Therefore, the bispectrum, which is defined at late time $\tau$, is independent of $\tau$ up to NLO as a consequence of the cancellation of $\ln(-\tau)$, meaning that the $\tau$-dependence of $H,\varepsilon_{i}$ in (3.1) with (3.2) indeed cancels the logarithm $\ln(-\tau)$ in $\mathscr{A}_{\text{NLO}}$. Furthermore, one can confirm that the bispectrum (3.11) is independent of the choice of the pivot time $\tau^{*}$ by expanding $H^{*},\varepsilon_{i}^{*}$ around yet another pivot time $\tau^{\circ}$, which gives (3.11) with the replacement $(H^{*},\varepsilon_{i}^{*},\tau^{*})\to(H^{\circ},\varepsilon_{i}^{\circ},\tau^{\circ})$.
3.2 Dominant contributions to non-gaussianity and its conserved definition